OpenStudy (anonymous):

Please help! Tina is planning to fence in an area for her dog. She wants to create a rectangular area of 1,600 square feet for her pet and can afford to purchase 160 feet of fence. ?

OpenStudy (zpupster):

LW = 1600 and L+L+W+W = 160 or 2L+2W = 160 or 160-2W = 2L L=(160-2W)/2 substitute this for L in the first equation and solve

OpenStudy (anonymous):

you have just saved my life. thank you.